Identify the solution(s) of the system of equations, if any.
-6x - 4y = 6
12y = -18x - 18
I cannot remember how to solve this problem. Can you help?

Found 2 solutions by ewatrrr, richwmiller:
Answer by ewatrrr(24785)

3(-6x - 4y = 6)
18x + 12y = -18
0 = 0 Infinite number of solutions (same line)

Answer by richwmiller(17219)

12y = -18x - 18
divide by -3
-4y=6x+6
which is the first equation so they are the same line with an infinite number of solutions